# Fine-Tuning Large Language Models to Translate: Will a Touch of Noisy Data in Misaligned Languages Suffice?

### v1

@@ 4.1 What does LLM SFT bring us?
Foundational language models become more robust and follow instructions better after being fine-tuned on task-oriented supervised data formulated as natural language text (Mishra et al., 2022; Sanh et al., 2022; Wei et al., 2022). We observe diverging trends in research on instruction tuning nowadays: 1) Many works attempt to scale up instruction data in terms of the number of tasks, languages, data size, and thus implicitly increasing training updates (Chung et al., 2022; Muennighoff et al., 2023; Wu et al., 2024b; Li et al., 2023a; Üstün et al., 2024; Zhang et al., 2024). 2) Another stream of papers, argue that instruction tuning mainly alters a base model’s response style but not content or knowledge—data quality and diversity outweigh quantity (Zhou et al., 2023; Mitchell et al., 2023; Lin et al., 2024; Chen et al., 2024a). This work is a continued exploration of the latter, focusing on the machine translation task. We verify the effect of size variations and include two new factors—language directions and quality—aiming to provide practical and cost-effective guidance on this matter.
@@
Specifically, language transfer has been demonstrated in smaller pre-trained models before LLMs (Wu and Dredze, 2019; Artetxe et al., 2020). For (sufficiently) multilingual models, training on certain languages might still benefit other languages at the test time (Choenni et al., 2023). In LLM instruction tuning, recent papers revealed cross-lingual transfer and improved robustness in unseen languages via multilingual instruction tuning with a small data sample (Chen et al., 2024b; Kew et al., 2023; Shaham et al., 2024). Furthermore, it has been claimed that even monolingual instruction tuning is sufficient to elicit multilingual responses in the correct languages with a key ingredient being the right learning rate (Chirkova and Nikoulina, 2024a, b). In relation to our experiments, language transfer to unseen languages might account for improved performance in language directions that are not directly fine-tuned on.
4.2 Using LLMs for translation
@@
In the field of machine translation, earlier works provided analysis of general-purpose prompting (Vilar et al., 2023; Agrawal et al., 2023; Zhang et al., 2023a) followed by a blossom of strategies focusing on specific aspects of the translation process (Sarti et al., 2023; Ghazvininejad et al., 2023; He et al., 2023; Moslem et al., 2023; Chen et al., 2023; Raunak et al., 2023). Nonetheless, as shown in our experimental results, few-shot prompting is not on par with using instruction-tuned models, illustrating the importance of further understanding the role of instruction tuning in translation tasks.
@@
In terms of fine-tuning LLMs for translation, previous works have explored a wide range of sub-tasks: disambiguation, low-resource, document-level, and adaptive translation, etc (Li et al., 2023b; Zhang et al., 2023b; Alves et al., 2023; Iyer et al., 2023; Mao and Yu, 2024; Wu et al., 2024a). These works focus on improving translation performance and specific applications. Recent research aims to enhance the translation capabilities of LLMs by incorporating human preference data (Jiao et al., 2023; Zeng et al., 2024; Zhu et al., 2024) or by extending the pre-training phase before fine-tuning (Xu et al., 2024a, b; Alves et al., 2024), yet these approaches require significantly more data or computing resources. The aim of this paper is not to pursue the state of the art but to investigate the opportunities of extending instruction-tuned LLMs’ translation capabilities in desirable compute-efficient scenarios. It is still worth noting that our investigation is orthogonal to previous works which employ relatively large monolingual and parallel data for continued pre-training.
@@ 5 Conclusion and Future Work
In this work, we conduct an in-depth analysis of fine-tuning LLMs for translation. We demonstrate that LLMs is capable of translating in multiple directions after being fine-tuned with minimal low-quality training data in a single direction. While this suggests pre-trained LLMs inherently possess multilingual translation capabilities which only need to be unlocked by aligning with the correct task format, we discover pitfalls and lessons in aligning LLMs; while LLMs make efforts to adjust to the translation task, they are good at imitating other patterns such as the noise in the parallel data. Future work could explore robust training methods that align LLMs with translation while minimizing the risk of overfitting to noisy data.
@@
Throughout our systematic study, we fine-tuned Llama-2 7B, which is a strong and feasible option when the work is carried out. This may impact the generalization of our findings to models with different capabilities or of different sizes are used.
@@
Our evaluation is English-centric, which is the condition of most LLM pre-training; the findings will be more comprehensive if future work can extend it to translation directions not involving English.
@@
Our research primarily explores how SFT enables LLM to translate, with the aim of uncovering data-efficient strategies in SFT and identifying associated pitfalls. Recent studies have demonstrated that translation capabilities can be further enhanced through techniques such as continual pre-training (Xu et al., 2024a; Alves et al., 2024) and preference learning (Xu et al., 2024b; Zhu et al., 2024). However, these methods require significantly more training resources, which may pose challenges when applied to large models.
@@
Throughout this work, we perform full parameter fine-tuning in SFT. It is worthwhile to explore alternative tuning methods and investigate whether they exhibit properties similar to those observed in full fine-tuning, especially when training involves small, single-direction, and noisy data. We leave this investigation for future work.
@@ Acknowledgments
Pinzhen Chen and Barry Haddow received funding from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee [grant number 10052546].
@@ Appendix D Model Performance with Unseen Languages
In Figure 9, we find similar patterns as the COMET score, where fine-tuning on unseen languages can elicit the model’s ability to translate from and to all seen languages. However, the translation performance on unseen languages themselves remains under-performed, suggesting that SFT primarily reveals the knowledge LLMs have possessed during pre-training.
@@ F.3 Evaluation packages
To obtain COMET scores, we use the Unbabel/wmt22-comet-da model888https://github.com/Unbabel/COMET, and for BLEU score evaluations, we use sacreBLEU999https://github.com/mjpost/sacrebleu. The signature from the sacreBLEU package is nrefs:1, case:mixed, eff:no, tok:13a, smooth:exp, version:2.0.0 for all language pairs , except for enzh and enjp, where we use tok:zh and tok:jp-mecab, respectively.
@@ F.4 Hardware specifications and runtime
Our experiments are conducted on a computing node with either 8 NVIDIA A100-40GB GPUs or 8 H100-80GB GPUs. Deepspeed101010https://github.com/microsoft/DeepSpeed with zero-stage 1 and mixed precision bfloat16 is used for performing SFT. Given the limited dataset size, typically fewer than 1024 samples, each SFT experiment can be completed within a mere 15 minutes using four H100 GPUs. However, given the necessity to evaluate the models across more than ten translation directions, the evaluation process may require up to four hours when performed on a single A100-40GB GPU.

### v2

@@ 4.1 What does LLM SFT bring us?
Foundational language models become more robust and follow instructions better after being fine-tuned on task-oriented supervised data formulated as natural language text (Mishra et al., 2022; Sanh et al., 2022; Wei et al., 2022). We observe diverging trends in research on instruction tuning nowadays: (1) Many works attempt to scale up instruction data in terms of the number of tasks, languages, data size, and thus implicitly increasing training updates (Chung et al., 2024; Muennighoff et al., 2023; Wu et al., 2024c; Li et al., 2023; Üstün et al., 2024; Zhang et al., 2024). (2) Another stream of papers, argue that instruction tuning mainly alters a base model’s response style but not content or knowledge—data quality and diversity outweigh quantity (Zhou et al., 2023; Mitchell et al., 2024; Lin et al., 2024; Chen et al., 2024a). This work is a continued exploration of the latter, focusing on the machine translation task. We verify the effect of size variations and include two new factors—language directions and quality—aiming to provide practical and cost-effective guidance on this matter.
@@
Specifically, language transfer has been demonstrated in smaller pre-trained models before LLMs (Wu and Dredze, 2019; Artetxe et al., 2020). For (sufficiently) multilingual models, training on certain languages might still benefit other languages at the test time (Choenni et al., 2023). In LLM instruction tuning, recent papers revealed cross-lingual transfer and improved robustness in unseen languages via multilingual instruction tuning with a small data sample (Chen et al., 2024c; Kew et al., 2023; Shaham et al., 2024). Furthermore, it has been claimed that even monolingual instruction tuning is sufficient to elicit multilingual responses in the correct languages with a key ingredient being the right learning rate (Chirkova and Nikoulina, 2024a, b). In relation to our experiments, language transfer to unseen languages might account for improved performance in language directions that are not directly fine-tuned.
4.2 How can we use LLMs for translation?
@@
In the field of machine translation, earlier works provided analysis of general-purpose prompting (Vilar et al., 2023; Agrawal et al., 2023; Zhang et al., 2023a) followed by a blossom of strategies focusing on specific aspects of the translation process (Sarti et al., 2023; Ghazvininejad et al., 2023; He et al., 2024; Moslem et al., 2023; Chen et al., 2024b; Raunak et al., 2023). Nonetheless, as shown in our experimental results, few-shot prompting is not on par with using instruction-tuned models, illustrating the importance of further understanding the role of instruction tuning in translation tasks.
@@
In terms of fine-tuning LLMs for translation, previous works have explored a wide range of sub-tasks: disambiguation, low-resource, document-level, and adaptive translation, etc (Li et al., 2024; Zhang et al., 2023b; Alves et al., 2023; Iyer et al., 2023; Mao and Yu, 2024; Wu et al., 2024b). These works focus on improving translation performance and specific applications. Stap et al. (2024) show that while fine-tuning improves translation quality, it can degrade certain key LLMs’ advantages, such as the contextualization ability on document-level input. Some recent research aims to enhance the translation capabilities of LLMs by incorporating human preference data (Jiao et al., 2023; Zeng et al., 2024; Zhu et al., 2024) or by extending the pre-training phase before fine-tuning (Xu et al., 2024a, b; Alves et al., 2024), yet these approaches require significantly more data or computing resources. The aim of this paper is not to pursue the state of the art but to investigate the opportunities of extending instruction-tuned LLMs’ translation capabilities in desirable compute-efficient scenarios. It is still worth noting that our investigation is orthogonal to previous works which employ relatively large monolingual and parallel data for continued pre-training.
@@ 5 Conclusion and Future Work
In this work, we conduct an in-depth analysis of fine-tuning LLMs for translation. We demonstrate that LLMs is capable of translating in multiple directions after being fine-tuned with minimal low-quality training data in a single direction. While this suggests pre-trained LLMs inherently possess multilingual translation capabilities which only need to be unlocked by aligning with the correct task format, we discover pitfalls and lessons in aligning LLMs; while LLMs make efforts to adjust to the translation task, they are good at imitating other patterns such as the noise in the parallel data. Future work could explore robust training methods that align LLMs with translation while minimizing the risk of overfitting to low-quality data.
@@
Throughout our systematic study, we fine-tuned Llama-2 7B, Llama-2 12B, and Mistral 7B. These are strong and feasible options when the work is carried out. It is important to verify the generalizability of our findings to models with different capabilities or of different sizes.
@@
Our evaluation is English-centric, which is the condition of most LLM pre-training. Findings will be more comprehensive if future work can extend it to translation directions not involving English.
@@
Our research primarily explores how SFT enables LLM to translate to uncover data-efficient strategies in SFT and identify associated pitfalls. Recent studies have demonstrated that translation capabilities can be further enhanced through techniques such as continual pre-training (Xu et al., 2024a; Alves et al., 2024) and preference learning (Xu et al., 2024b; Zhu et al., 2024). However, these methods require significantly more training resources, which may pose challenges when applied to large models.
@@
Throughout this work, we perform SFT with full-parameter updates. It is worthwhile to explore parameter-efficient methods which bring in heavier regularization to understand whether they exhibit patterns similar to those observed in our work.
@@ Acknowledgments
We sincerely thank the reviewers of this work for their constructive and insightful feedback.
Pinzhen Chen and Barry Haddow received funding from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee [grant number 10052546]. Miaoran Zhang received funding from the DFG (German Research Foundation) under project 232722074, SFB 1102. We thank EIT and IDT High Performance Computing Center for providing computational resources for this project.
@@ Appendix A Model Performance with Varying Training Sample Sizes
While we primarily report the results with Llama-2 7B in our experiments, we hypothesize that state-of-the-art LLMs are largely homogeneous in terms of language distribution and inherent translation capability making our findings applicable to other LLMs. To support this hypothesis, we conduct fine-tuning experiments with Mistral 7B and Llama-2 13B using varying data sizes: 32, 1024, and 70K. As shown in Figure 8, the general trend is quite similar to the Llama-2 7B case: fine-tuning with 32 examples results in competitive performance, matching or surpassing general-purpose instruction-tuned models. Furthermore, increasing the number of training examples leads to diminishing returns.
@@ Appendix D Model Performance with Unseen Languages
In Figure 10, we find similar patterns as the COMET score, where fine-tuning on unseen languages can elicit the model’s ability to translate from and to all seen languages. However, the translation performance on unseen languages themselves remains subpar, suggesting that SFT primarily reveals the knowledge LLMs have possessed during pre-training.
@@ F.3 Evaluation packages
To obtain COMET scores, we use Unbabel/wmt22-comet-da999https://github.com/Unbabel/COMET and for BLEU scores, we use sacreBLEU101010https://github.com/mjpost/sacrebleu (Post, 2018). The signature from the sacreBLEU package is nrefs:1, case:mixed, eff:no, tok:13a, smooth:exp, version:2.0.0 for all language pairs, except for tokenization for enzh and enjp, where we use tok:zh and tok:jp-mecab, respectively.
@@ F.4 Hardware specifications and runtime
Our experiments are conducted on a computing node with either 8 NVIDIA A100-40GB GPUs or 8 H100-80GB GPUs. DeepSpeed111111https://github.com/microsoft/DeepSpeed with zero-stage 1 and mixed precision bfloat16 is used for performing SFT. Given the limited dataset size, typically fewer than 1024 samples, each SFT experiment can be completed within a mere 15 minutes using four H100 GPUs. However, given the necessity to evaluate the models across more than ten translation directions, the evaluation process may require up to four hours when performed on a single A100-40GB GPU.